A multiple choice exam has ten questions and each question had 4 possible answers. In each case only 1 of the 4 possibilities is correct. Billy fills out an answer sheet by simply guessing without even reading the questions. What is the probability that he will get 100% on the exam? What is the probability that he will get 90% on the exam? What is the probability that he will get at leat an 80% on the exam?
Answer by stanbon(75887)

A multiple choice exam has ten questions and each question had 4 possible answers. In each case only 1 of the 4 possibilities is correct.
------------------------------------------------
Billy fills out an answer sheet by simply guessing without even reading the questions.
----
Binomial with n = 10 and p = 1/4
----
What is the probability that he will get 100% on the exam?
P(x = 10) = (1/4)^10 = 9.5367x10^-7
------------------------------------------
What is the probability that he will get 90% on the exam?
P(x = 9) = 10C9*0.25^9*0.75 = 2.8610x10^-5
------------------------------------------
What is the probability that he will get at least an 80% on the exam?
P(x = 8) = 10C8*0.25^8*0.75^2 = 3.8624x10^-4
----------------------------------------------
Cheers,
Stan H.
===========